A fresh produce distributor uses 890 non-returnable packing crates a month, which it purchases at a cost of $6 each. The manager has assigned an annual holding cost of 35 percent of the purchase price per crate. Ordering cost is $35 per order. Currently the manager orders 890 crates at a time. How much could the firm save annually in ordering and holding costs by using the EOQ?
Answer #1

EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST)

ANNUAL HOLDING COST = (EOQ / 2) * HOLDING COST

ANNUAL ORDERING COST = (DEMAND / EOQ) * ORDERING COST

EOQ = SQRT(2 * 10680 * 35 / 2.1) = 596.66

ANNUAL HOLDING COST = (596.66 / 2) * 2.1 = 626.49

ANNUAL ORDERING COST = (10680 / 596.66) * 35 = 626.49

TOTAL COST = AHC + AOC = 626.49 + 626.49 = 1252.98

TOTAL COST OF ORDERING 890 AT A TIME

ANNUAL HOLDING COST = (890 / 2) * 2.1 = 934.5

ANNUAL ORDERING COST = (10680 / 890) * 35 = 420

TOTAL COST = AHC + AOC = 934.5 + 420 = 1354.5

SAVINGS =1354.5 - 1252.98 = $ 100.52

Do give a positive rating if you think this answer helped